A hiker, who weighs 708 N, is strolling through the woods and crosses a small horizontal bridge. The bridge is uniform, weighs 3040 N, and rests on two concrete supports, one on each end. He stops 1/3 of the way along the bridge. What is the magnitude of the force that a concrete support exerts on the bridge (a) at the near end and (b) at the far end?
Answer #1


Net torque about an axis passing through near support

Torque = rxF

N2*L = (L/3)*708 + (3040)(L/2)

N2 = (708/3)+(3040/2)

force at far end N2 = 1756 N


Net torque about an axis passing through far support

N1*L = (2L/3)(708) +(3040)(L/2)

N1 = (2*708/3) +(3040/2)

force at near end N1 =1992 N